THE PURE ROTATIONAL SPECTRA OF CrS (X^{5}Π_{r}): CONTINUED STUDIES OF THE 3d TRANSITION METAL SULFIDES

\underline{R. L. PULLIAM}, A. J. HIGGINS, and L. M. ZIURYS, Department of Chemistry, Department of Astronomy, Steward Observatory, University of Arizona, Tucson, AZ 85721.

The pure rotational spectrum of CrS has been measured in the ground X^{5}Π_{r} state using gas-phase direct absorption methods in the frequency range of 280-405 GHz. The molecule was created by the reaction of chromium vapor, sublimed in a Broida oven, with hydrogen sulfide gas. All five spin components were observed in eleven rotational transitions, and lambda-doubling splitting were resolved in the Ω = 0, 1 and 2 ladders. The data were fit with a Hunds case(a) Hamiltonian and rotational, spin-orbit, spin-spin, and lambda-doubling constants have been established. A bond length of 2.0781 Å was derived from the data. Using the lambda-doubling parameters, the energy difference between the ground X^{5}Π and excited ^{5}Σ state was estimated to be 1400 cm ^{−1}.